### v1

# Embodied LLM Agents Learn to Cooperate in Organized Teams

Abstract
Large Language Models (LLMs) have emerged as integral tools for reasoning, planning, and decision-making, drawing upon their extensive world knowledge and proficiency in language-related tasks. LLMs thus hold tremendous potential for natural language interaction within multi-agent systems to foster cooperation. However, LLM agents tend to over-report and comply with any instruction, which may result in information redundancy and confusion in multi-agent cooperation. Inspired by human organizations, this paper introduces a framework that imposes prompt-based organization structures on LLM agents to mitigate these problems. Through a series of experiments with embodied LLM agents and human-agent collaboration, our results highlight the impact of designated leadership on team efficiency, shedding light on the leadership qualities displayed by LLM agents and their spontaneous cooperative behaviors. Further, we harness the potential of LLMs to propose enhanced organizational prompts, via a Criticize-Reflect process, resulting in novel organization structures that reduce communication costs and enhance team efficiency.
1 Introduction
Modern intelligent systems, such as autonomous vehicle networks and swarms of drones, often involve complex decision-making processes where multiple agents or components must collaborate seamlessly to achieve specific objectives Wang et al. (2020); Vinyals et al. (2019); Zhang et al. (2019a); Wang et al. (2021). In these systems, communication among the various agents is pivotal, as it dictates the flow of information, coordination of tasks, and overall system performance Zhang et al. (2019b); Guo et al. (2023); Foerster et al. (2016); Das et al. (2019). Agents in traditional multi-agent systems often have to communicate in pre-specified ways, such as exchanging gradients, sharing data, state observations and actions, etc Kim et al. (2020); Lin et al. (2021); Foerster et al. (2016). The emergence of large language models (LLMs) makes it possible for AI agents to communicate and cooperate using natural language, bringing enormous flexibility and potential for more nuanced and human-understandable interactions Park et al. (2023); Hong et al. (2023); Mandi et al. (2023); Chen et al. (2023a).
Despite the flexibility of LLMs, integrating them into practical multi-agent systems remains a challenge. While LLMs are trained and finetuned for text generation and instruction-following, they are not necessarily tailored to multi-agent cooperation. Modern LLMs are prone to over-reporting and obeying instructions, as a by-product of RLHF finetuning (Bai et al., 2022), and they can ignore critical information (Liu et al., 2023a) or be distracted by irrelevant information (Shi et al., 2023), especially when the context is long (see Figure 1 for examples). While recent studies involving agent-based LLMs have demonstrated they are capable of solving problems through multi-agent collaboration Li et al. (2023b); Zhang et al. (2023c); Mandi et al. (2023), it is worth noting that such collaborations often follow predefined patterns designed using heuristics to channel the behavior of the models productively Li et al. (2023b). Creating systems that support free-flowing interaction between LLMs in a way that could potentially scale to include human collaborators is still an open problem.
This paper investigates the collaborative potential of LLM agents working in teams. Drawing on prior studies in human collaboration from cognitive and economic perspectives, there is potential for organizations to be redesigned to more effectively manage the limited attention span within teams, as suggested by Simon et al. (1971), and mitigate individual limitations and enhance overall team performance, as highlighted by Van Zandt (1999) and Vélez et al. (2023). Specifically, we study two research questions. First, what role do organizational structures play in multi-LLM-agent systems? Second, how can we optimize these organizational structures to support efficient multi-agent coordination? By leveraging AutoGen Wu et al. (2023), a generic multi-agent conversation framework, we develop a framework for studying how to best organize embodied LLM agents to communicate and collaborate in physical/simulated non-text environments Zhang et al. (2023c). Our framework offers the flexibility to prompt and organize LLM agents into various team structures, facilitating versatile inter-agent communication. It also serves as a testbed to empirically evaluate the traditional ideas proposed in the organization theory literature.
Our initial experiments in this setting reveal that uncoordinated LLM agents often send redundant and repetitive messages and interrupt others’ actions, leading to chaos (see Fig. 1 and Appendix D). To remedy these issues, we explore organizational structures that allow multiple LLM agents to collaborate and complete a common task efficiently.
The first organizational structure we explore is a hierarchy, a classic object of study in organizational theory March and Simon (1958); Radner (1993); Chisholm (1992); Bolton and Dewatripont (1994); Garicano (2000); Dodds et al. (2003). With a designated leader, LLM agents work more efficiently and collaboratively. For the example of a three-agent team, imposing a leader improves efficiency by up to 30% with almost no extra communication cost (up to 3%), consistent with findings for human organizations Dodds et al. (2003). This also holds true in five-agent cases. Further, LLM agents demonstrated the potential to elect their own leader and adjust leadership dynamically via communication. With proper organizations, LLM agents exhibit a variety of cooperative behaviors that mimic humans. For example, agents can provide constructive suggestions and seek help from others; they can also execute appropriate interactions for a hierarchy such as reporting back on task progress; see Figures 6 and 7 and Appendix C. We also tested human-agent collaboration, and observe that, unsurprisingly, human leaders are much better at coordinating a team of agents when compared to AI agents.
Figure 1: Example of disorganized communication and interruption, without a designated leader. In a team of three GPT-4 agents, two agents engaged in unnecessary communication and made disordered decisions, causing a delay due to the lack of a predefined organization. We identified many more examples including conflicting messages and repetitive communications, see Appendix D.
In addition to testing existing organizational structures, we explore the use of LLMs to improve the organization prompt. To this end, we develop a Criticize-Reflect framework, adopting a dual LLM architecture, to reflect on the team performance and generate improved and novel organization prompts. Through this iterative process, our LLM agents spontaneously form novel, effective team structures, leading to reduced communication costs and improved efficiency; see Figures 8 and 9.
2 Related Works
2.1 LLM Agents
As powerful LLMs inherit abundant world knowledge and also general reasoning ability, there are increasing efforts to deploy LLMs as the reasoning core for decision-making to build human-like autonomous agents (Sun et al., 2023; Zhu et al., 2023; Hao et al., 2023a). This requires observations of the RL environment to be translated into natural language in a way that is easier for language models to process. The reasoning of the language models also needs to be turned into a viable action for execution. Popular prompting techniques for doing so include ReAct (Yao et al., 2022) and Reflexion (Shinn et al., 2023). Other methods that involve fine-tuning the language models have also been explored (Hao et al., 2023b). In addition, various techniques have been proposed to mitigate the biases and constraints of language models, including chain-of-thought reasoning (Wei et al., 2022), external tools (Shen et al., 2023; Patil et al., 2023), external documents (Wang et al., 2023) and skill libraries (Zhu et al., 2023).
2.2 Multi-Agent Cooperation
Multi-agent cooperation has been extensively studied for decades under various topics such as communication efficiency, planning, leadership, and team dynamics using different platforms (Lowe et al., 2017; Samvelyan et al., 2019; Resnick et al., 2018; Puig et al., 2021) (see recent surveys for detail (Oroojlooy and Hajinezhad, 2023; Zhang et al., 2021; Gronauer and Diepold, 2022)). Previous works mainly focused on communication through continuous vectors (Das et al., 2019) or discrete symbols (Lowe et al., 2017; Jaques et al., 2019).
Recent work (Xu et al., 2023; Zhang et al., 2023d; Wu et al., 2023; Li et al., 2023a) showed that multiple LLM agents or human-agent teams can improve upon single LLM in solving pure text-based tasks, such as creative writing, reasoning, and code generation. Liu et al. (2023b), Hong et al. (2023) and Zheng et al. (2023) further explored agent selection or role assignment to improve the performance.
LLMs have also been applied to multi-agent cooperation for embodied tasks (Agashe et al., 2023; Mandi et al., 2023; Park et al., 2023; Chen et al., 2023a). Besides, Zhang et al. (2023b) proposed an intention inference framework to enhance the cooperation of LLM agents without explicit communication. Li et al. (2023b) investigated LLM-agents collaboration for Theory of Mind inferences tasks with a broadcast-only communication protocol and homogeneous policies. Zhang et al. (2023c) studied embodied multi-agent cooperation in the two-agent and the one-human-one-agent settings. Chen et al. (2023b) explored different fixed communication structures for multi-LLM-robots. These initial explorations are also limited to fixed team structures and are not optimized for communication efficiency. In contrast, our work explores the impact of deploying and optimizing organizational structures, allowing $\geq 3$ agents in a team, for efficient multi-agent communication and cooperation.
2.3 Prompt Optimization
Language models are sensitive to prompts. The format of the prompt can have a substantial influence on performance (Gao et al., 2020; Wei et al., 2022; Zhou et al., 2022a; Shi et al., 2023; Zou et al., 2023; Qi et al., 2023). Various research efforts have aimed at improving performance through prompt optimization. Typical approaches include heuristic search using language models’ knowledge (Gao et al., 2020; Shin et al., 2020), first-order methods like soft prompt tuning (Lester et al., 2021), and prefix tuning (Li and Liang, 2021). In this work, we focus on obtaining an interpretable prompt in the form of natural language, drawing on insights from Yang et al. (2023), Zhou et al. (2022b), and Pryzant et al. (2023).
3 Method
3.1 Architecture and Multi-Agent Communication
We adopt the embodied LLM-agent architecture proposed by Zhang et al. (2023c) and expand it to enable organized teams of $\geq 3$ agents to communicate, plan, and act in physical/simulated environments. Figure 2 illustrates our architecture. Borrow insights from Zhang et al. (2023c), we adopt four standard modules: Configurator, Perception Module, Memory Module, and Execution Module. They are responsible for configuring the agents, translating environmental observations into text, storing & retrieving historical information, and executing actions, respectively (Fig. 2(a)).
Figure 2: Multi-LLM-agent architecture. (a) The modules of an LLM agent and the composition of prompts. (b) There are two phases in one time step: Communication phase and Action phase. In the communication phase, the agents take turns communicating by broadcasting or selecting receivers to send distinct messages. The agents can also choose to keep silent. Comm is short for Communication; PO is short for Partial Observation.
Figure 3: Criticize-Reflect architecture for improving organizational structure. The red agent represents the leader in an organization. The Critic evaluates the trajectories and analyzes the agents’ performance. Together with the external costs from the environment, the Coordinator proposes a new organization prompt to improve the team efficiency.
Previous works focused on two-agent cooperation, in which case the communication can be simply treated as an extra action Mandi et al. (2023); Zhang et al. (2023c). In contrast, we aim to enable three or more agents to work in a team and cooperate through organized communication. Thus we design the architecture with several features that facilitate organized multi-agent communication (Figure 2(b)):
  • We impose an organizational structure for the agent team via prompting, i.e., including a textual description as part of the prompt for each round of communication.
  • LLM agents keep alternating between two phases during their task: the communication phase and the action phase. The standalone communication phase supports richer team structures and flexible communication patterns.
  • During communication, agents take turns to communicate. An agent can choose to broadcast a message, select one recipient for a message, choose multiple recipients and send them distinct messages, or remain silent. Agents keep their own history of communication and can respond to messages from previous communications.
3.2 Criticize-Reflect Method for Improving Organizational Structure
We leverage powerful LLMs to optimize the organization prompt, borrowing insights from (Yang et al., 2023). To do so, we introduce a dual-LLM framework to allow the multi-LLM-agent system to ponder and improve the organizational structure. Figure 3 illustrates the architecture of our framework. It consists of two LLMs:
  • LLM critic: Inspired by the Actor-Critic method of reinforcement learning Konda and Tsitsiklis (1999), we introduce an LLM critic to evaluate the team’s performance based on verbal feedback. The team critic takes as input the dialogue and action history of one episode. Then, the critic analyzes the input and reasons to extract and summarize the key steps that are believed to influence the performance in the episode. Also, the critic provides a textual evaluation of agents’ behaviors and the ranking of their leadership. Note that the critic in our method is different from Zhang et al. (2023a) where the critic is a centralized controller offering suggestions to each agent. See the prompts in Appendix A.
  • LLM coordinator: The LLM coordinator takes as input the outputs of the LLM critic as well as cost metrics of previous episodes from the environment. It reflects on these data and generates thoughts based on the analysis of the past episodes. The coordinator then generates three distinct new prompts and chooses the best out of them. For more details please refer to Appendix A.
For each new organization prompt, we run for one episode following the new organizational structure and then return the dialogue and action history to the critic. By alternating between criticizing and reflecting on current structure, the framework discovers more effective, novel organizational structures with self-improvement.
3.3 Environment Setup
We chose VirtualHome-Social Puig et al. (2018, 2021) as the environment and extended it to support multi-LLM-agent communication and interaction. In this environment, agents are humanoid helpers in a virtual home doing housekeeping, where the tasks include Prepare afternoon tea, Wash dishes, Prepare a meal, Put groceries, Set up a dinner table, etc. For instance, in Figure 1, the agents cooperate to prepare afternoon tea by searching for and transporting task-specific items (chocolate, juice, wine, etc.) to a target location (the coffee table). The environment generates symbolic observations of the objects in the home and their relations. Each agent only observes the objects in the open containers located in her room and teammates in the same room, but she can walk to another room to explore. Any agent can communicate with any other agent, not subject to a range limit.
Each episode starts from an initial state where agents are randomly located in the environment and all containers are closed. The episode terminates when the task is fully completed. To evaluate the team’s efficiency we measure the number of time steps taken to task completion, and we report the average number of tokens communicated between agents per step. In our experiment, each run initializes with an independently randomized state to obtain the mean and a confidence interval. We adopt GPT-4, GPT-3.5-turbo (Ouyang et al., 2022), and Llama2-70B (Touvron et al., 2023) as LLMs in our agents. The temperature is set as 0.8, the maximum number of output tokens is 256, and the number of completion choices to generate is 1.
4 Main Results
Figure 4: Organized teams with a designated leader achieve higher efficiency. (a,b) Comparison between the case of disorganized agents, the case where a leader is appointed, the case where agents choose their own leader dynamically, and the case where a human player replaces an agent to be the leader. Note that GPT-3.5-turbo doesn’t support leadership election. (c,d) Comparing leadership quality for GPT-3.5-turbo vs. GPT-4. The confidence intervals of Human as the leader group are calculated over 3 seeds while others are over 20 seeds.
Figure 5: Examples of communication messages when there is a designated leader. Left: messages from lead agents; Right: messages from non-lead agents. GPT-4 (upper), GPT-3.5-turbo (center), and Llama2-70B (lower) demonstrated different communication styles.
4.1 A Designated Leader Enhances Performance
We first studied the effect of organizational structures and leadership on LLM agents. For benchmarking, we experimented with disorganized LLM agents without providing any organization prompt. In this case, agents still communicate with one another and work to complete the overall task. However, we discovered frequent occasions where agents send redundant, repetitive messages and interfere with one another. See Figure 1 for an illustration and see Appendix D for more examples. Numeric metrics are reported in Appendix Table 1.
[…]
The proposed multi-LLM-agent architecture is also human-friendly to support human-AI collaboration. In the experiment, we ask a human player to replace the leader in the team of 3 GPT-4 agents. We recruit three human players to conduct the experiments. Figure 4(a, b) demonstrates that human leadership achieved better task completion time and improved communication efficiency compared with GPT-4 as the leader. Please find more examples of dialogues between the human leader and LLM agents in Appendix F.
4.2 Leadership and Open Communication Matters
LLM agents have different levels of leadership. In the team with a mixture of GPT-4 and GPT-3.5-turbo agents, appointing GPT-4 as the leader increases the team efficiency higher than if GPT-3.5-turbo is the leader (Figure 4(c,d)). We ran this experiment on teams of three agents and five agents, respectively. In both scenarios, the task completion time and communication cost are reduced when GPT-4 acts as the leader. This finding implies different levels of leadership between these LLMs.
We also observed that encouraging constructive feedback to the leader agent helped performance. Motivated by successful human organizations, we tried to promote open communications among LLM agents by adding an additional prompt that "If the leader’s instructions are not right, you can correct the leader". Figure 4(c, d) illustrates the results. Interestingly, this modification improves the team’s overall efficiency and reduces the time to task completion when the team is made up of 3$\times$×GPT-4 ($t(38)=0.87,p=0.14$). In contrast, the same modification lowers the team efficiency when GPT-3.5-turbo agents try to correct the leader ($t(38)=0.27,p=0.40$). In both experiments, the communication cost increases. We present more details about these behaviors in Appendix G.
4.3 Emergence of Cooperative Behaviors
We delved into the behaviors of LLM agents in an organized team to investigate how organization prompts influence agents’ communication and decisions. Analysis of their dialogue history revealed that agents demonstrated a variety of cooperative behaviors, such as reporting, correction, task allocation, and asking for help (see Figure 6 for an example dialogue).
One may argue that these types of behaviors could also emerge due to the nature of LLMs, even without a pre-specified team structure. Thus we performed a quantitative analysis to study the impact of an organization prompt on these behaviors. We followed a three-step process:
  • (1)
    We defined three major categories of human cooperative behaviors: (i) Information sharing: agents influence others by offering new information, either actively or by being asked. Reporting to the leader, sharing new observations, and answering questions belong to this category. (ii) Leadership & assistance: agents, especially the leader if there is one, can influence others by changing their plans. The behaviors include task allocation, correction, and asking for help. (iii) Request for guidance: agents actively request new information or plans for their own decision making.
  • (2)
    We developed a standalone prompt-based GPT-4- classifier to analyze each piece of dialogue. The classifier decides whether to label the dialogue with any subset of the aforementioned labels. The classifier has an accuracy of 91.67% when tested on 20 human-labeled dialogue samples with 60 labels (see Appendix A for the prompt and Appendix H for the test samples).
  • (3)
    We use the classifier to label messages generated by our teams of agents, and report the percentages of messages with cooperative behaviors. Note that a message may have multiple labels.
Figure 7 reports the results and illustrates the behavior patterns for different LLM agents. The results support several observations. Even in a disorganized team, LLM agents love to tell others what to do. Leadership & assistance accounts for around $>$ 50% of all the behaviors (Figure 7(a)). However, other than telling others what to do, agents in the disorganized team do not show much cooperative behaviours, for example they would request for guidance in $<10\%$ of the dialogues.
In contrast, when the team has a hierarchical organization the lead LLM agent would presume a dominant role and give orders to others (amount to $>60\%$ of their communication), while other members tend to follow and give fewer orders compared with the disorganized case. (Figure 7(b, c)). In such a team, agents tend to share and ask for more information, especially for the follower agents in the team.
Figure 6: Examples of cooperative behaviors in a dialogue. Agent_3 leads the team (3$\times$×GPT-4 agents). The agents emerge three types of cooperative behaviors: information sharing, leadership & assistance, and request for guidance.
Figure 7: Emergent cooperative behaviors of LLM agents. We analyzed the communication log of the mixture team (1$\times$×GPT-4+2$\times$×GPT-3.5-turbo) and asked another GPT-4 to annotate agent’s cooperative behaviors. (a) Behavior of disorganized agents. (b) Behavior of a team led by a GPT-4 agent. (c) Behavior of a team led by a GPT-3.5-turbo agent.
4.4 Novel Organizational Structures
Having evaluated the merits of different kinds of structures, we let the LLM propose novel organizational structures and iteratively refine the organization prompts using the Criticize-Reflect method discussed in Section 3 (see also Figure 3).
Figure 8(a) visualizes the reflection process. The system was initialized with a basic organization prompt, i.e., “Agent_1 as the leader to coordinate the task". As the Reflection process moves forward, the Coordinator generates a sequence of evolving organization prompts, picking up key words like “hierarchical" and “dynamic" that imply more complex team structures.
[…]
The Critic analyzes the records of action and dialogue, and performance metrics from the most recent episode. It provides evaluation for the full team’s trajectory, feedback to individual agents and their rankings. See below for an example of the Critic output:
[…]
In addition, it is worth mentioning that LLMs are able to generate highly complex prompts that imply novel organizational structures that are rarely seen in human societies. We illustrate the communication patterns as team structures in Figure 9 together with the three novel structures proposed by Criticize-Reflect: (c) chain, (d) dual-leader, and (e) dynamic structures, which are the best structures of the three settings in Figure 8(b, c) respectively.
[…]
Figure 8: The reflection and improvement process for finding novel organizational structures. (a) The experiment was done using the 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo team. The organization prompt evolves during the iterations, and takes on additional keywords such as "central", "hierarchical", and "dynamic". (b) The confidence intervals are calculated over 20 seeds.
Figure 9: Communication patterns and the corresponding organizational prompts for different team structures. (a) Team without organization prompts. (b) Team with a leader. (c) A team in the chain structure. (d) A dual-leader team. (e) A team with a dynamic leadership. (c, d, e) are proposed by LLM via Reflection. Red-robot nodes mark the lead agents, and other nodes are the followers. Edges mark the accumulated communication cost between the two nodes (darker edge means higher token costs).
5 Conclusion
We developed a novel multi-LLM-agent architecture to facilitate communication and organize the embodied agent teams for enhanced cooperation. Moreover, we proposed the Criticize-Reflect architecture based on LLMs to generate more efficient organizational prompts. Extensive experiments with various group settings and organizational structures demonstrate that a hierarchically-organized team with a designated/elected leader has superior team efficiency, which can be further improved by Criticize-Reflect.
The current work is performed in a single environment and lacks human evaluation. Future work shall expand the evaluation to a broader set of environments, allowing human evaluation and more complex human-AI collaboration.
Impact Statement
This research studies the integration of prompt-based organizational structures to teams of LLM agents, contributing to more efficient and coherent multi-agent interactions. These findings have the potential to greatly influence the deployment of more effective and autonomous multi-agent systems in various fields, including robotics, virtual assistants, etc. For example, the study has potential applications in disaster response scenarios, where efficient multi-agent coordination is crucial.
Appendix A Prompt Templates
We list the prompts of Actor, Communicator, Critic, and Coordinator as follows.
Actor and the Communicator. ORGANIZATION_INSTRUCTION is the placeholder for the organization instruction prompt, either manually designed or automatically generated. The environment will provide text descriptions for the current GOAL, PROGRESS, and AVAILABLE_ACTIONS. We include the latest $12$ sent and received messages as DIALOGUE_HISTORY, and the latest $20$ steps of actions as ACTION_HISTORY.
Critic. We provide the full trajectory as the input to TRAJECTORIES. Additionally, ORGANIZATION_INSTRUCTION and GOAL of the current task and organization are also provided as an additional context.
Coordinator. In “Instruction examples”, we include the basic setting (goal, organization structure instruction), the communication cost, the number of steps taken, as well as the summarized information generated by the Critic (leadership ranking, problems, summary of the trajectory) for the Coordinator.
Classifier. We feed the messages to the GPT-4 classifier and get the labels. The rubrics are manually written after investigating the communication logs.
Appendix B Additional Results
B.1 Complete list of basic experimental results
We present the full results of various group settings and organization instructions in Appendix Table 1. Here, we also include the results of 1$\times$×GPT-4+2$\times$×Llama2-70B. Surprisingly, GPT-4 exhibits poorer leadership than Llama2-70B in this case. The communication costs for the teams containing Llama2-70B are much higher than those containing GPT-3.5-turbo.
Table 1: Performance for different organization instructions. When there are two different kinds of LLMs in the group, Agent_1 is GPT-4, and Agent_2 is the other type of LLM.
Group setting Organization instruction Time Communication cost 3$\times$×GPT-4 None 57.75 $\pm$13.09 67.03 $\pm$9.68 3$\times$×GPT-4 Agent 1 is the leader to coordinate the task. 54.70 $\pm$8.92 54.73 $\pm$8.89 3$\times$×GPT-4 Agent 1 is the leader to coordinate the task. If the leader’s instructions are not right, you can correct the leader. 50.70$\pm$13.92 63.49$\pm$8.61 3$\times$×GPT-4 Elect a new leader every 10 steps to coordinate the task. … After the election, the other agents should follow the leader’s instructions. 49.20$\pm$9.97 135.03$\pm$20.45 3$\times$×GPT-3.5-turbo None 102.95$\pm$21.88 53.73$\pm$6.04 3$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. 92.90$\pm$14.70 59.87$\pm$6.33 3$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. If the leader’s instructions are not right, you can correct the leader. 94.20$\pm$16.22 60.53$\pm$3.66 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo None 81.10$\pm$18.35 54.00$\pm$5.06 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. 73.30$\pm$16.12 55.82$\pm$6.57 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. If the leader’s instructions are not right, you can correct the leader. 85.67$\pm$14.52 61.57$\pm$0.55 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo Agent 2 is the leader to coordinate the task. 75.65$\pm$15.43 58.39$\pm$8.11 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo Agent 2 is the leader to coordinate the task. If the leader’s instructions are not right, you can correct the leader. 72.33$\pm$6.60 74.21$\pm$7.73 1$\times$×GPT-4+2$\times$×Llama2-70B None 77.00$\pm$2.94 119.48$\pm$1.28 1$\times$×GPT-4+2$\times$×Llama2-70B Agent 1 is the leader to coordinate the task. 83.67$\pm$10.96 135.22$\pm$16.39 1$\times$×GPT-4+2$\times$×Llama2-70B Agent 2 is the leader to coordinate the task. 76.00$\pm$5.72 142.24$\pm$11.85 2$\times$×GPT-4+3$\times$×GPT-3.5-turbo None 42.67$\pm$4.03 98.03$\pm$9.86 2$\times$×GPT-4+3$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. 39.67$\pm$9.46 94.73$\pm$4.01 2$\times$×GPT-4+3$\times$×GPT-3.5-turbo Agent 2 is the leader to coordinate the task. 48.50$\pm$9.50 96.53$\pm$2.51
B.2 Across Task Generalizability
We conduct experiments across the tasks to test the generalizability of the prompt “dynamic leadership” (Figure 9(e)) found using Criticize-Reflect architecture on the Prepare_Afternoon_Tea task and report the performance in Figure 10; see Section 4.4 for the complete setting and discussions.
Figure 10: The organized team structure with a designated leader and the novel structure proposed by Criticize-Reflect architecture generalized to different tasks. The prompt for dynamic leadership is proposed by Criticize-Reflect architecture on the Prepare_Afternoon_Tea task shown in Figure 8(a). The experiment was done using the 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo team over two seeds for each task. (a, b) Hard tasks (read_book, put_dishwasher_hard, prepare_food) with typical numbers of steps to accomplish the tasks $>60$. (c, d) Easy tasks ( put_dishwasher_easy, put_fridge, setup_table) with typical numbers of steps to accomplish the tasks $<60$.
Appendix C Emergent Cooperative Behaviors in an Organization
By investigating the messages between agents, we mainly observe the following cooperative behaviors, as summarized in Table 2.
Table 2: Typical cooperative behaviors.
Type Description Example Sharing information An agent shares her observations to others, reports her task-related progress to others, or responds to other agents’ requests (Ex 1.)“I’m in the bathroom. There’s an unchecked $<$bathroomcabinet$>$ (190).”
(Ex 2.) “I’ll check the cabinet in the bedroom ”
Giving orders An agent gives orders to others, either by directly giving a command or by a polite request “I still need to find $<$pudding$>$ (371). Can you help me search the bedroom for the remaining item?” Asking for information An agent asks other agents about their location, task progress, or other information (Ex 1.) “Where are you now?”
(Ex 2.) “Any updates from the kitchen?”
(Ex 3.) “Do we know the location of the coffeetable?”
Exchanging information An agent shares one agent’s information to another agent Agent 3 $\to$ Agent 1: “ Found cupcake and juice in bedroom, plus a wine. ”;
Agent 1 $\to$ Agent 2: “Agent3 found a wine, cupcake, and juice in the bedroom.”
Asking for orders An agent asks what she needs to do or whether she can help others “I’m in the kitchen with the dishwasher, stove, microwave unchecked. Need me to check these or something else?” Correction An agent corrects the plan of others (Ex 1.) Agent 2 $\to$ Agent 3: “After I place the pudding on the coffeetable, let’s split up. You check the bathroom, I’ll check the bedroom.”; Agent 3 $\to$ Agent 2: “You don’t need to search the bedroom, the juice (380) is there. ”
(Ex 2.) “There are no more wine or juice in the kitchen cabinets. We need to come up with another plan.”
Appendix D Ineffective Communication
There are also cases in which language model agents fail to communicate efficiently. From the messages between agents, we summarize the typical categories in Table 3.
Table 3: Typical ineffective communication behaviors.
Type Description Example Duplicated message The same message is said more than once Agent 1 $\to$ Agent 3: “Check kitchencabinet or dishwasher for wine, I’ll check the stove.”,
Agent 2 $\to$ Agent 3: “Agent 1 will check the stove for wine.”
Repeated command The same command is said more than once Agent 3 $\to$ Agent 1: “Explore the bathroom for the second wine.”
Agent 2 $\to$ Agent 1: “Go to the bathroom and look for the wine. ”
Conflicting command An agent gives another agent an order that conflicts with her current task Agent 1 $\to$ Agent 3: “Find $<$juice$>$ and check $<$bathroomcabinet$>$ (190)”
Agent 2 $\to$ Agent 3: “Check the kitchen containers for the last wine.”
Conflicting plans An agent claims to commit to two conflicting plans Agent 3 $\to$ Agent 1: “I found the pudding and will check the dishwasher, stove, and microwave for the second wine.” Agent 3 $\to$ Agent 2: “I will continue searching for the second wine in the bedroom and bathroom.” Improper delegation An agent asks another agent to do her own task Agent 1 $\to$ Agent 2: “Continue checking the kitchen cabinets for remaining items.”
Agent 2 $\to$ Agent 3: “Please continue checking the other kitchen cabinets for the remaining items. ”
Ignoring requests An agent ignores other agents’ questions Agent 2 $\to$ Agent 3: “I haven’t found any of the remaining items in the kitchen. Have you found any of the required items in the living room?”
Agent 3 $\to$ Agent 2: “I haven’t explored the bathroom yet.”
Appendix E Examples of Election
In Figure 11, the agents vote to elect a new leader. We can observe behaviors such as nominations for themselves and other agents, voting, and consensus achievement. We find that the agents are not power-seeking and may give up leadership early. The agents prefer to vote for others instead of nominating themselves (5 times more during the whole task). The elected leader also does not plan to keep the position but to nominate others for the next round. Also, the agents’ standpoint can be easily influenced by others. The agents do not debate much to win the election but reach a consensus soon. For example, Agent_1 gives up running for herself but votes for Agent_2 because of Agent_3’s support. Furthermore, sometimes nominations and votes are determined by hallucinations. For example, at step 2, Agent_2 nominates Agent_1 as he was the first one to propose a search strategy. However, based on the previous dialogues, Agent_1 has not proposed any strategy yet.
Figure 11: Examples of the election of a new leader. It takes two steps to vote and negotiate to determine the new leader in this case. Note that Agent_3 chooses not to send a message as the election is done and no more information to be shared for now. All the messages in the figure are broadcasts.
Appendix F Examples of Human-AI Collaboration
We conducted experiments involving a team consisting of one human player and two GPT-4 agents, with the human player acting as the leader. Figure 12 illustrates the remarkable collaboration between humans and AI.
Figure 12: Examples of human-AI collaboration when the human player leads two GPT-4 agents (Agent_2&3).
Appendix G Examples of Correction
Due to hallucination and the limit of the dialogue history buffer, the leader may forget what has happened and give wrong orders. When the prompt encourages the agents to correct the leader when necessary by adding If the leader’s instructions are not right, you can correct the leader, some correction behaviors appear, as shown in Figure 13.
In the first example, the leader Agent_1 gives an unnecessary and repetitious instruction. Then Agent_2 corrects the leader to avoid time wasting. In the second example, the leader Agent_1 may have hallucinations and cannot remember what Agent_3 is holding clearly (cupcake and wine in the message while juice and wine in the thoughts). Therefore, Agent_3 clarifies that she is not holding the cupcake and wine and shares her next plan with the leader.
Figure 13: Examples of correction dialogues and the corresponding thoughts. The prompt includes If the leader’s instructions are not right, you can correct the leader.
Appendix H Examples of Cooperative Behaviors Classification by Humans and GPT-4
We ask a human evaluator and GPT-4 to label the dialogues into three different behavior categories shown in Table 4.
Table 4: Dialogues and behavior labels given by humans and GPT-4. Label 1: information sharing; Label 2: leadership & assistance; Label 3: Request for guidance.
Dialogue Human GPT-4 Label 1* Label 2 Label 3 Label 1 Label 2 Label 3 Hey, where are you? Please let me know your location so that I can assign you a task. 0 1 1 0 1 1 I’m currently in the bedroom where I found an unchecked cabinet. Please explore the livingroom and start checking for the required items. 1 1 0 1 1 0 I’m currently in the bedroom where I found an unchecked cabinet. I haven’t found any of the required items yet. 1 0 0 1 0 0 Check the cabinet in the bedroom. I’ll check the one in the bathroom. 1 1 0 0 1 0 I haven’t found any of the required items yet. Did you find any in the kitchen? 1 0 1 1 0 1 I haven’t found any of the required items yet. Have you found any in the bathroom? 1 0 1 1 0 1 Agent 3, I found a pudding in one of the kitchen cabinets. Please continue checking the other kitchen cabinets for the remaining items. 1 1 0 1 1 0 I haven’t found any of the required items yet. Did you find any other required items in the kitchen? 1 0 1 1 0 1 I haven’t found any of the remaining items yet. 1 0 0 1 0 0 I’m in the living room searching for the remaining items. 1 0 0 1 0 0 Okay, I will keep checking the kitchen cabinets for the remaining items. 1 0 0 1 0 0 I’m currently in the living room searching for the remaining items. 1 0 0 1 0 0 I haven’t found any of the remaining items yet. 1 0 0 1 0 0 I’m still searching the living room. Have you found any of the required items? 1 0 1 1 0 1 Put items on the table once you find them. Keep searching your current areas. 0 1 0 0 1 0 I haven’t found any of the remaining items in the kitchen. Have you found any of the required items in the living room? 1 0 1 1 0 1 I haven’t explored the bathroom yet. Have you found any of the required items in the living room? 1 0 1 1 0 1 I have explored the bedroom and found wine, cupcake, and juice. I still need to find pudding. Can you help me search the bedroom for the remaining item? 1 1 1 1 1 0 I found the wine, cupcake, and juice in the bedroom. Agent 1 wants me to put them on the coffee table and then check the bathroom cabinet. 1 0 0 1 1 0 Agent 1 wants us to check if there’s another wine in the kitchen. 1 0 0 0 1 0
Appendix I Examples of New Prompts after Reflection
We list more prompts generated by the Criticize-Reflect architecture in Figure 14.
Figure 14: Examples of Prompts generated via Reflection. The first row is generated with the Critic, while the second row is without the Critic, where the new prompts are relatively vague. Note that there is no Agent Z in the team.